OpenStudy (anonymous):

Solve, using the substitution method. 2x + 4y = 16 x – 7 = – 4y

OpenStudy (anonymous):

1. x+2y=6 or x=6-2y 3x-2y=2 3(6-2y)-2y=2 18-6y-2y=2 18-8y=2 -8y=-18+2 -8y=-16 y=-16/-8 y=2 ans. x+2*2=6 x+4=6 x=6-4=2 ans. Proof: 3*2-2*2=2 6-4=2 2=2 I leave the rest for you to practice with. 3. 4x+y=7 2x+5y=-1 5. 3x+2y=-2 6x-y=6 Solve the system using the elimination method. Show all your steps. 7. -3x+3y=3 3x+y=9 add. --------------------------------- 4y=12 y=12/4 y=3 ans. 3x+3=9 3x=9-3 3x=6 x=6/3 x=2 ans. Proof: -3*2+3*3=3 -6+9=3 3=3 I'll leave the rest for practice. 9. -5x+12y=20 x-2y=-6 11. 3x+2y=1 4x+ 6y= 7

OpenStudy (anonymous):

A. (0, -7) B. (0, 7/4) C. (9, -1/2) D. (17, -5/2)

OpenStudy (luigi0210):

Sub in the second into the first, so: x – 7 = – 4y becomes: x=7-4y Now plug it into the first equation: \[\LARGE 2(7-4y) + 4y = 16\] \[\LARGE 14-8y+4y=16\] Combine like terms: \[\LARGE -4y=2\] Now divide: \[\LARGE y=-\frac{1}{2} \] So now we have the y-value: \[\LARGE (x, -\frac{1}{2})\] Can you find x?

OpenStudy (anonymous):

yes, thank you

OpenStudy (luigi0210):

You're welcome!